# Middleware for LLMs: Tools Are Instrumental for Language Agents in Complex Environments

### v1

@@ 5.2 Main Results
As shown in Tables 1 and 2, equipping LLMs with customized tools leads to significant improvement over previous standards, almost doubling or tripling the performance under multiple metrics. Specifically, API docs prompting can only feed the schema information to the LLM due to the vast amount of database content. As a result, it fails catastrophically on examples that require database content to compose the SQL query. In contrast, Fuxi equips the agent with tools to actively navigate the database to collect relevant information for composing a SQL query. As a result, Fuxi significantly closes the gap between performance on questions requiring database content and questions not requiring it when using GPT-4 (i.e., $45.1$% vs. $38.3$%). Additionally, we notice that Fuxi minimizes the gap between with and without oracle knowledge from $15.5$% to $4.0$% using GPT-4 and $11.7$% to $3.3$% using GPT-3.5-turbo. Finally, StructGPT demonstrates a similar trend to API docs prompting because its tools do not provide any information about the database content. For KBQA, Fuxi demonstrates uniformly superior performance across different question types and significantly outperforms Pangu with both GPT-3.5-turbo and GPT-4. In particular, when equipped with GPT-4, Fuxi + decoupled generation outperforms Pangu by $32.2$% in F1. As for the other two baselines, KB-Binder and StructGPT, both fail miserably on our challenging setting. On the one hand, KB-Binder only retrieves relations within two hops from the entities for grounding. However, most questions in KBQA-Agent involve more than two relations. As a result, many of its drafted programs are unable to ground, which explains its low VA. On the other hand, StructGPT is heavily limited by its constrained toolset and cannot handle complex questions in KBQA-Agent. Therefore, StructGPT frequently refuses to provide an answer (as revealed by its low VA) due to insufficient information. The strong performance of Fuxi underscores that tools are instrumental for language agents in complex environments.
@@ 5.3 Experiments with Open-Source LLMs
To gain a more thorough insight, we also include experiments with four open-source LLMs ( Figure 3). Our findings indicate that Llama2 models generally underperform compared to other LLMs, aligning with trends observed in other LLM leaderboards, such as Chatbot Arena Zheng et al. (2023). Specifically, we find Llama2 models struggle with even generating grammatical tool use following our instruction. On the other hand, Mistral and Mixtral demonstrate much better performance than Llama2. In particular, Mixtral represents an advanced mixture-of-experts model that has demonstrated superior performance and even surpasses GPT-3.5-turbo on Chatbot Arena Zheng et al. (2023). However, different from answering open-ended questions featured in Chatbot Arena, properly engaging with the complex environment demands the language agent to produce more precise actions that strictly conform to the task specification. There is still a performance gap between Mixtral and GPT-3.5-turbo in terms of predicting valid actions over intricate environments. Compared to GPT-3.5-turbo, Mixtral tends to output invalid actions more frequently. This also explains why decoupled generation, where the output space is strictly constrained to a list of valid actions, helps weaker models more. With Fuxi + decoupled generation, using Mistral can almost match the best baseline performance with GPT-3.5-turbo, and using Mixtral can even match the best baseline with GPT-4. While stronger models like GPT-4 can effectively recover the mistake via error feedback, weaker models tend to benefit more from decoupled generation.
@@ 5.4 Tools as A Middleware Layer
To deepen our understanding of the integral roles of tools in aiding LLMs in accessing complex environments (i.e., KB triples and database rows in our setup), we conduct further analysis by comparing Fuxi with prompting baselines with different amounts of data items directly sampled from the environment (Figure 4). For the KB, we sample $10$, $50$, $100$, and $200$ triples from Freebase based on the three-hop neighborhood of each entity in a question. These triples are the top-ranked ones using a sentence-BERT retriever Reimers and Gurevych (2019) based on their similarity with the input question. We prompt the LLM directly with these sampled triples and request it to generate an answer to the given question. Given the extensive size of Freebase, accurately representing the environment with a mere subset of samples proves to be exceedingly difficult. Consequently, both GPT-3.5 Turbo and GPT-4 consistently yield an F1 score close to 00. For the database, we similarly augment API docs prompting with $1$, $5$, and $10$ sampled rows for each table and evaluate on $100$ random questions from Bird that require accessing database content. Additionally, we also augment Fuxi with the same sampled rows in the database setting. We observe that including more database rows initially boosts baseline performance but eventually decreases it. With Fuxi, prompting the LLM with sampled rows yields minimal gain, and the standard setting without sampled rows already significantly outperforms all baselines. These results further confirm that the LLM, when augmented with tools, can effectively engage with complex environments, flexibly gathering the necessary information on demand and bypassing the limitations on the amount of data it can handle (e.g., around $200$ triples or $10$ rows per table).
@@ 6 Conclusion
In this paper, we aim to address the compelling question we posed: how effectively can LLMs handle complex environments with the aid of tools? We investigate this through evaluations in two exemplary environments: KBs and databases. While we achieve notable results in these environments, it is important to acknowledge that implementing customized tools for KBs and databases presents fewer challenges compared to environments without a straightforward query interface, such as a webpage or a physical environment. In future work, we plan to extend Fuxi across a broader range of environments, aiming to fully realize the potential of language agents in complex environments through the integration of customized middleware tools.
Furthermore, the tools developed in this study are soley grounded in our experience. Despite this, our results already demonstrate the significant potential of augmenting LLMs with customized tools in complex environments, aligning with the primary objective of this paper. Nonetheless, to enhance performance further, adopting a more principled strategy for tool design is essential.
Acknowledgements
We would like to thank colleagues from THU KEG and the OSU NLP group for their thoughtful comments. In addition, we are grateful to the authors of Bird for kindly providing the original predictions of the baselines based on API Docs prompting. This research was supported by a sponsored award from Cisco Research.

### v2

@@ 5.2 Main Results
As shown in Tables 1 and 2, equipping LLMs with customized tools leads to significant improvement over previous standards, almost doubling or tripling the performance under multiple metrics. Specifically, API docs prompting can only feed the schema information to the LLM due to the vast amount of database content. As a result, it fails catastrophically on examples that require database content to compose the SQL query. In contrast, Middleware equips the agent with tools to actively navigate the database to collect relevant information for composing a SQL query. As a result, Middleware significantly closes the gap between performance on questions requiring database content and questions not requiring it when using GPT-4 (i.e., $45.1$% vs. $38.3$%). Additionally, we notice that Middleware minimizes the gap between with and without oracle knowledge from $15.5$% to $4.0$% using GPT-4 and $11.7$% to $3.3$% using GPT-3.5-turbo. Finally, StructGPT demonstrates a similar trend to API docs prompting because its tools do not provide any information about the database content. For KBQA, Middleware demonstrates uniformly superior performance across different question types and significantly outperforms Pangu with both GPT-3.5-turbo and GPT-4. In particular, when equipped with GPT-4, Middleware + decoupled generation outperforms Pangu by $32.2$% in F1. As for the other two baselines, KB-Binder and StructGPT, both fail miserably on our challenging setting. On the one hand, KB-Binder only retrieves relations within two hops from the entities for grounding. However, most questions in KBQA-Agent involve more than two relations. As a result, many of its drafted programs are unable to ground, which explains its low VA. On the other hand, StructGPT is heavily limited by its constrained toolset and cannot handle complex questions in KBQA-Agent. Therefore, StructGPT frequently refuses to provide an answer (as revealed by its low VA) due to insufficient information. The strong performance of Middleware underscores that tools are instrumental for language agents in complex environments.
Due to the space limit, we provide additional results in Appendix D.
@@ 5.3 Experiments with Open-Source LLMs
To gain a more thorough insight, we also include experiments with four open-source LLMs ( Figure 3). Our findings indicate that Llama2 models generally underperform compared to other LLMs, aligning with trends observed in other LLM leaderboards, such as Chatbot Arena Zheng et al. (2023). Specifically, we find Llama2 models struggle with even generating grammatical tool use following our instruction. On the other hand, Mistral and Mixtral demonstrate much better performance than Llama2. In particular, Mixtral represents an advanced mixture-of-experts model that has demonstrated superior performance and even surpasses GPT-3.5-turbo on Chatbot Arena Zheng et al. (2023). However, different from answering open-ended questions, properly engaging with the complex environment demands the language agent to produce more precise actions that strictly conform to the task specification. There is still a gap between Mixtral and GPT-3.5-turbo in terms of predicting valid actions over complex environments. Compared to GPT-3.5-turbo, Mixtral tends to output invalid actions more frequently. This also explains why decoupled generation, where the output space is strictly constrained to a list of valid actions, helps weaker models more. With Middleware + decoupled generation, using Mistral can almost match the best baseline performance with GPT-3.5-turbo, and using Mixtral can even match the best baseline with GPT-4. While stronger models like GPT-4 can effectively recover the mistake via error feedback, weaker models tend to benefit more from decoupled generation.
@@ 5.4 Tools as A Middleware Layer
To deepen our understanding of the integral roles of tools in aiding LLMs in accessing complex environments (i.e., KB triples and database rows in our setup), we conduct further analysis by comparing Middleware with prompting baselines with different amounts of data items directly sampled from the environment (Figure 4). For the KB, we sample $10$, $50$, $100$, and $200$ triples from Freebase based on the three-hop neighborhood of each entity in a question. These triples are the top-ranked ones using a sentence-BERT retriever Reimers and Gurevych (2019) based on their similarity with the input question. We prompt the LLM directly with these sampled triples and request it to generate an answer to the given question. Given the extensive size of Freebase, accurately representing the environment with a mere subset of samples proves to be exceedingly difficult. Consequently, both GPT-3.5 Turbo and GPT-4 consistently yield an F1 score close to 00. For the database, we similarly augment API docs prompting with $1$, $5$, and $10$ sampled rows for each table and evaluate on $100$ random questions from Bird that require accessing database content. Additionally, we also augment Middleware with the same sampled rows in the database setting. We observe that including more database rows initially boosts baseline performance but eventually decreases it. With Middleware, prompting the LLM with sampled rows yields minimal gain, and the standard setting without sampled rows already significantly outperforms all baselines. These results further confirm that the LLM, when augmented with tools, can effectively engage with complex environments, flexibly gathering the necessary information on demand and bypassing the limitations on the amount of data it can handle (e.g., around $200$ triples or $10$ rows per table).
@@ 6 Conclusion
In this paper, we aim to address the compelling question we posed: how effectively can LLMs handle complex environments with the aid of tools? We investigate this through evaluations in two exemplary environments: KBs and databases. While we achieve notable results in these environments, it is important to acknowledge that implementing customized tools for KBs and databases presents fewer challenges compared to environments without a straightforward query interface, such as a webpage or a physical environment. In future work, we plan to extend Middleware across a broader range of environments, aiming to fully realize the potential of language agents in complex environments through the integration of customized middleware tools.
Furthermore, the tools developed in this study are soley grounded in our experience. Despite this, our results already demonstrate the significant potential of augmenting LLMs with customized tools in complex environments, aligning with the primary objective of this paper. Nonetheless, to enhance performance further, adopting a more principled strategy for tool design is essential. Additionally, investigating autonomous tool-making methods Wang et al. (2024) in complex environments presents a promising direction for future research.
@@ Appendix C Prompts
Appendix D Additional Results
Avg Input Tokens Avg Time (s) Avg Rounds Perceivable Rows EX (100 questions)
Middleware (error feedback) 14 257.214257.214\,257.214 257.2 $15.5$ $7.1$infty $39.0$
API Docs Prompt $1296.8$ $2.5$ $1.0$ 00 $12.0$
API Docs Prompt (10 rows) $4635.5$ $2.9$ $1.0$ $10$ $22.0$
API Docs Prompt (20 rows) $8459.4$ $2.9$ $1.0$ $20$ $9.0$
(a) Table 1
Avg Input Tokens Avg Time (s) Avg Rounds Perceivable Triples F1
Middleware (error feedback) 25 140.725140.725\,140.725 140.7 $22.1$ $9.8$infty $55.1$
Middleware (decoupled generation) 23 331.823331.823\,331.823 331.8 $20.9$ $9.3$infty $59.3$
Pangu-Chat $4675.7$ $67.3$ $4.5$infty $27.1$
Direct Prompt (200 triples) $7654.5$ $3.1$ $1.0$ $200$ $0.8$
(b) Table 2
Table D.2: The average running time and input tokens for different methods. Note that, no existing research addresses the challenge of handling extensive DB content, limiting existing methods to perceiving only a small number of rows. This is a critical gap in the literature that we aim to fill.
Setting gpt-3.5-turbo-0613 gpt-4-0613
w/ all $14.0$ $38.0$
w/o find_columns_containing_value $14.0$ $38.0$
w/o find_columns_containing_value_fuzzy $12.0$ $34.0$
w/o get_distinct_values $10.0$ $32.0$
w/o is_value_in_column $14.0$ $36.0$
w/o all four $8.0$ $22.0$
Table D.3: Ablation study on different tools used in our DB tasks.
D.1 Comparing with DIN-SQL and DAIL-SQL
In addition to API Docs Prompt, we also compare with two strong baselines with their source code available: DIN-SQL Pourreza and Rafiei (2023) and DAIL-SQL Gao et al. (2024) from Bird’s leaderboard. Note that in their original submission, they only evaluated under the oracle knowledge setting. To ensure a fair comparison with Middleware under the same without oracle knowledge setting, we adapted their source code to evaluate without oracle knowledge with minimal changes. Specifically, some modules in their pipeline design require several in-context demonstrations, which we also preserved. Consequently, we use 6-shot for DIN-SQL and 7-shot for DAIL-SQL, following their original codebase. In particular, we randomly sample $100$ questions from Bird’s dev set and evaluate them on these questions. The results is presented in Table D.4.
D.2 Efficiency Analysis
We further look into the efficiency (e.g., avg. tokens and avg. running time) of different methods to gain a deeper understanding. For DBs, we compare Middleware with the most straightforward single-round baseline API Docs Prompt. For KBs, we compare Middleware with Pangu, as all other baselines significantly underperform. The concrete results are presented in Table D.2.
D.3 Ablation for Tools
For KBQA, an ablation experiment for tools would be trivial as each tool deterministically contributes to the final KB queries. For example, removing get_relations or get_neighbors would yield 0% performance on all questions, while removing count would result in 0% performance on counting questions. As a result, we only do the ablation experiment for text-to-SQL parsing. Specifically, we first identify the top-4 commonly used tools based on GPT-4’s predictions: find_columns_containing_value, find_columns_containing_value_fuzzy, get_distinct_values, and is_value_in_column. On 50 out of the previous 100 questions (due to the budget concern), we show the results in Table D.3. An intriguing observation is the robustness exhibited when removing only one tool, owing to the redundancy built into our tool designs. Nevertheless, when all four tools were removed, the performance experienced a substantial decline.
gpt-3.5-turbo-0613 gpt-4-0613
Middleware (0-shot) $16.0$ $39.0$
API Doc Prompt (0-shot) $6.0$ $12.0$
DIN-SQL (6-shot) - $22.0$
DAIL-SQL (7-shot) $8.0$ $11.0$
Table D.4: Results of more baselines on Bird w/o oracle knowledge.